# UDC: A Unified Neural Divide-and-Conquer Framework for Large-Scale Combinatorial Optimization Problems

### v1

@@ 3.2 Training Method: Divide-Conquer-Reunion (DCR)
Both the dividing and conquering stages in neural divide-and-conquer methods can be modeled as MDP [12]. The reward for the conquering policy is derived from the objective function of the sub-solutions, while the reward for the dividing policies is calculated based on the objective function of the final merged solution. Existing methods face challenges of instability when training both policies simultaneously with RL [20]. Therefore, they generally employ a separate training scheme, pre-training a conquering policy on specific datasets before training the dividing policy [13]. However, this separate training process not only requires additional problem-specific dataset designs [13] but also undermines the collaborative optimization of the dividing and conquering policies (further discussed in Section 5.1).
In this paper, we imply that the instability of unified training is related to the sub-optimal sub-problems decomposition. As shown in Figure 2, sub-problems generated by the initial solution will probably not match any continuous segments in the optimal solution so there may be errors emerging at the connection (i.e., linking region) of these sub-solutions. The naive solving procedure of neural divide-and-conquer methods ignores the negative impact of these solution errors, thus obtaining biased rewards for the dividing policy. To tackle this drawback, we propose the Divide-Conquer-Reunion (DCR) process to eliminate the impact as much as possible. DCR designs an additional Reunion step to treat the connection between the original two adjacent subproblems as a new sub-problem and perform another conquering stage on $\boldsymbol{x}_{1}$ with new sub-problem decompositions. The Reunion step can be formulated as rolling the original start point $x_{1,0}$ along the solution $\boldsymbol{x}_{1,0}$ by $l$ times (i.e., to $x_{1,l}$) and conquering again. To ensure equal focus on the two adjacent sub-problems, we set $l=\frac{n}{2}$ in training all the DCR-enabled UDC models. DCR provides a better estimation for the reward of dividing policies, thereby increasing the stability and convergence speed in training.
Therefore, the unified training scheme of UDC consists of a dividing policy training stage and two consecutive conquering policy training stages (rolling the starting point of the sub-problem by $\frac{n}{2}$ in the second conquering stage). UDC adopts the REINFORCE algorithm [43] in training both the dividing policy (with Loss $\mathcal{L}_{d}$) and the conquering policy (with Loss $\mathcal{L}_{c1}$ for the first conquering stage and $\mathcal{L}_{c2}$ for the second). The baseline in training dividing policy is the average reward over $\alpha$ initial solutions $\boldsymbol{x}^{i}_{0},i\in\{1,\ldots,\alpha\}$ and the baseline for conquering policy is calculated over $\beta$ sampled sub-solutions. The gradients of loss functions on a single instance $\mathcal{G}$ are calculated as follows:
@@
$\nabla\mathcal{L}_{c1}(\mathcal{G})$ $=\frac{1}{\alpha\beta\lfloor\frac{N}{n}\rfloor}\sum_{c=1}^{\alpha\lfloor\frac{N}{n}\rfloor}\sum_{i=1}^{\beta}\Big{[}\big{(}f(\boldsymbol{s}^{1,i}_{c},\mathcal{G}_{c}^{0})-\frac{1}{\beta}\sum_{j=1}^{\beta}f(\boldsymbol{s}^{1,j}_{c},\mathcal{G}_{c}^{0})\big{)}\nabla\ \log\pi_{c}(\boldsymbol{s}^{1,j}_{c}|\mathcal{G}_{c}^{0},\Omega_{c}^{0},\theta)\Big{]},$
$\nabla\mathcal{L}_{c2}(\mathcal{G})$ $=\frac{1}{\alpha\beta\lfloor\frac{N}{n}\rfloor}\sum_{c=1}^{\alpha\lfloor\frac{N}{n}\rfloor}\sum_{i=1}^{\beta}\Big{[}\big{(}f(\boldsymbol{s}^{2,i}_{c},\mathcal{G}_{c}^{1})-\frac{1}{\beta}\sum_{j=1}^{\beta}f(\boldsymbol{s}^{2,j}_{c},\mathcal{G}_{c}^{1})\big{)}\nabla\ \log\pi_{c}(\boldsymbol{s}^{2,j}_{c}|\mathcal{G}_{c}^{1},\Omega_{c}^{1},\theta)\Big{]},$
@@ 3.3 Application: Applicability in General CO Problems
This subsection discusses the applicability of UDC to various CO problems. Most existing neural divide-and-conquer methods rely on problem-specific designs, so they are only applicable to limited CO problems. UDC divides the original CO problems into their sub-problems with similar constraints, and uses no heuristics algorithms in the divide-and-conquer process, thus having the potential to be applied to general CO problems which satisfy the following three conditions:
@@
For the second condition, the proposed UDC framework leverages autoregressive solution generation procedures for both solutions and sub-solutions. Therefore, the proposed UDC is disabled on complex CO problems whose solution cannot be guaranteed to be legal through an autoregressive solution process (e.g., TSPTW). For the third condition, on certain CO problems such as the MIS problem on dense graphs or job scheduling problems, the solution of sub-problems has already been uniquely determined, so the conquering stages become ineffective.
@@ 4 Experiment
To verify the applicability of UDC in the general CO problem, we evaluate the UDC on 10 combinatorial optimization problems that satisfy the conditions proposed in Section 3.3, including TSP, CVRP, 0-1 Knapsack Problem (KP), Maximum Independent Set (MIS), ATSP, Orienteering Problem (OP), PCTSP, Stochastic PCTSP (SPCTSP), Open VRP (OVRP), and min-max multi-agent TSP (min-max mTSP). Detailed formulations of these problems are provided in the Appendix B.
Implementation By implementing lightweight networks for dividing policies, UDC can be directly trained on large-scale CO instances (e.g., TSP500 and TSP1,000). To learn more scale-independent features, we follow the varying-size training scheme [44, 45] in training and train one UDC model for each CO problem. The Adam optimizer is generally used for all involved CO problems and the detailed hyperparameters for each CO problem are provided in Appendix D.1. Most training tasks can be completed within 10 days on an NVIDIA Tesla V100S GPU.
Baselines We compare the proposed UDC to advanced heuristic algorithms and existing neural solvers for large-scale CO problems, including classical solvers (Concord [46], LKH [47], EA4OP [48], OR-Tools [49]), single-stage SL-based sub-path solvers (BQ [14], LEHD [8]), single-stage RL-based constructive solvers (AM [50], POMO[5], ELG [6], ICAM [16], MDAM [51], et al.), heatmap-based solvers (DIMES [35], DIFUSCO [39], et al.) and neural divide-and-conquer methods (GLOP [13], TAM [12], H-TSP[20], et al.). The implementation details and settings of all the baseline methods are described in Appendix G.
Table 3: Experiment results on 500-node to 2,000-node OP and PCTSP and SPCTSP. The 500-scale, 100-scale, and 2,000-scale problems contain 128, 128, and 16 instances, respectively. The overall best performance is in bold and the best learning-based method is marked by shade.
@@
Methods Obj. [^2] OP aims to maximize the objective function, which PCTSP and SPCTSP aim to minimize the objective function Gap Time Obj. [^2] OP aims to maximize the objective function, which PCTSP and SPCTSP aim to minimize the objective function Gap Time Obj. [^2] OP aims to maximize the objective function, which PCTSP and SPCTSP aim to minimize the objective function Gap Time
@@
Results on 500-node to 2,000-node Instances. We examine the performance of the model on large-scale CO problems ranging from 500-node to 2,000-node. Both single-stage SL-based constructive methods [16] and single-stage sub-path solvers [8] cannot be directly trained on this scale, so we report the results of models trained on the maximum available scale for these baselines. For UDC, we provide not only the greedy results (UDC-$\boldsymbol{x}_{2}$) but also the results after conducting more conquering stages (i.e., 50 and 250 stages, labeled as UDC-$\boldsymbol{x}_{50}$, UDC-$\boldsymbol{x}_{250}$). In the following conquering stages after a greedy solution, the starting point of fragment decomposition is no longer fixed at $\frac{n}{2}$ but is selected from Uniform$(0,n)$. This procedure is inspired by the random reconstruction (RRC) method in LEHD [8]. For better performances, we sample $\alpha=50$ initial solutions $\boldsymbol{x}_{0}$ for TSP and CVRP. The comparison results of TSP and CVRP are shown in Table 2 and the results of OP, PCTSP, and SPCTSP are exhibited in Tabel 3, UDC demonstrates the general applicability, exhibiting consistent outstanding performances in all the involved CO problems. After multiple conquering stages, UDC can achieve significantly better results.
@@
UDC variants demonstrate the best results in PCTSP, SPCTSP, CVRP1,000, and CVRP2,000, together with the best learning-based method in most involved CO problems. The UDC-$\boldsymbol{x}_{250}$($\alpha$=1) showcases extensive efficiency in OP, PCTSP, and PCTSP. Compared to sub-path solvers BQ [14] and LEHD [8], UDC variants exhibit more competitive results on TSP1,000, TSP2,000, CVRP, and OP. RL-based constructive solver ICAM [16] with 8 augmentation [5] is outstanding performance on TSP500, TSP1,000, and CVRP500 (i.e., relatively smaller scale) but UDC variants demonstrates better effect and efficiency in other test sets. Moreover, compared to existing neural divide-and-conquer methods GLOP [13], H-TSP [20], and TAM(LKH3) [12], UDC achieves significant advantages in performance with no heuristic design, which preliminarily implies the significance of a unified training scheme.
@@
Results on Benchmark Datasets. We also evaluate the proposed UDC on benchmark datasets TSPLib [52] and CVRPLib [53, 54]. We adopt the instances with 500-node to 5,000-node in TSPLib, CVRP Set-X [53], and very large-scale CVRP dataset Set-XXL as test sets. Instances in these datasets often with special distributions and can better represent real-world applications. As shown in Table 4, UDC variant UDC-$\boldsymbol{x}_{250}$ exhibits competitive benchmark results, demonstrating the best learning-based result on CVRPLib. The proposed UDC can also demonstrate superiority on the other 5 involved CO problems, very large-scale instances, and CO instances with different distributions. Results of these experiments are shown in Appendix D.2, Appendix D.4, and AppendixD.3, respectively.
@@ 5 Discussion
The experimental results have preliminarily proven that the proposed UDC method achieves excellent results on a wide range of large-scale CO problems with outstanding efficiency. In this section, we conduct ablation experiments to verify the necessity of components in the training processes of UDC. Appendix E further includes the ablation experiments on the testing procedure of UDC, including the ablation on the number of conquering stages $r$ and the number of sampled solutions $\alpha$.
@@ 5.1 Unified Training Scheme versus Separate Training Scheme
For neural divide-and-conquer methods, the unified training process is easier to implement than separate training, which does not require a special design for pre-training conquering policies. Moreover, unified training can also avoid the antagonism between the dividing and conquering policies. In this section, we pre-train a conquering policy (i.e., ICAM) on uniform TSP100 data and then train two ablation variants of the original UDC (separate training scheme): one variant uses the pre-trained conquering policy in the subsequent joint training (i.e., Pre-train + Joint Training, similar to H-TSP [20]), and the other variant only train the dividing policy afterward (i.e., Pre-train + Train Dividing, similar to TAM [12] & GLOP [13]). According to the training curves of UDC and the two variants in Figure 3, the unified training scheme in UDC demonstrates superiority, while the Pre-train step leads the training into local optima, which significantly harms the convergence of UDC.
@@
Figure 3: Training curves for ablation study.
Table 5: Results on various conquering policies.
@@
5.2 Ablation study
@@
In Appendix E.3, we conduct more ablation studies to demonstrate the necessity of other components of the proposed UDC, including the Reunion step in the DCR training method and the coordinate transformation in the conquering stage.
@@ 6 Conclusion, Limitation, and Future Work
This paper develops a training method DCR for neural divide-and-conquer methods by additionally considering the negative impact of sub-optimal dividing. By using DCR, this paper enables the superior unified training scheme and proposes a novel unified neural divide-and-conquer framework UDC. UDC exhibits not only outstanding performances but also extensive applicability, achieving significant advantages in 10 representative large-scale CO problems.
Limitation and Future Work. Although UDC has achieved outstanding performance improvements, we believe that UDC can achieve better results through designing better loss functions. In the future, we will focus on a more suitable loss for the UDC framework to further enhance the solution quality.
@@
In the Appendix, we provide a literature review A, detailed definitions of CO problems & sub-CO problems B, details of our proposed UDC method C, additional experimental results D, additional ablation study E, and visualizations on TSP1,000, CVRP1,000 (i.e., X-n1001-k43 in CVRPLib), and OP1,000 instances F.

### v2

@@ 3.2 Training Method: Divide-Conquer-Reunion (DCR)
Both the dividing and conquering stages in neural divide-and-conquer methods can be modeled as MDPs [12] (shown in Appendix C.7). The reward for the conquering policy is derived from the objective function of the sub-solutions, while the reward for the dividing policies is calculated based on the objective function of the final merged solution. Existing neural divide-and-conquer methods cannot train both policies simultaneously with RL [20]. Therefore, they generally employ a separate training scheme, pre-training a conquering policy on specific datasets before training the dividing policy [13]. However, this separate training process not only requires additional problem-specific dataset designs [13] but also undermines the collaborative optimization of the dividing and conquering policies (further discussed in Section 5.1).
In this paper, we imply that considering the negative impact of sub-optimal sub-problems decomposition is essential to implement the unified training of both policies. As shown in Figure 2, sub-problems generated by the initial solution will probably not match any continuous segments in the optimal solution so there may be errors emerging at the connection part (i.e., linking region) of these sub-solutions. The naive solving procedure of neural divide-and-conquer methods ignores the negative impact of these errors, thus obtaining biased rewards for the dividing policy. To tackle this drawback, we propose the Divide-Conquer-Reunion (DCR) process to eliminate the impact as much as possible. DCR designs an additional Reunion step to treat the connection between the original two adjacent sub-problems as a new sub-problem and perform another conquering stage on $\boldsymbol{x}_{1}$ with new sub-problem decompositions. The Reunion step can be formulated as rolling the original start point $x_{1,0}$ along the solution $\boldsymbol{x}_{1,0}$ by $l$ times (i.e., to $x_{1,l}$) and conquering again. To ensure equal focus on the adjacent two sub-problems, we set $l=\frac{n}{2}$ in training all the DCR-enabled UDC models. DCR provides a better estimation for the reward of dividing policies, thereby increasing the stability and convergence speed in unified training.
UDC adopts the REINFORCE algorithm [43] in training both the dividing policy (with Loss $\mathcal{L}_{d}$) and the conquering policy (with Loss $\mathcal{L}_{c1}$ for the Conquer step and $\mathcal{L}_{c2}$ for the Reunion step). The baseline in training dividing policy is the average reward over $\alpha$ initial solutions $\boldsymbol{x}^{i}_{0},i\in\{1,\ldots,\alpha\}$ and the baseline for conquering policy is calculated over $\beta$ sampled sub-solutions. The gradients of loss functions on a single instance $\mathcal{G}$ are calculated as follows:
@@
$\nabla\mathcal{L}_{c1}(\mathcal{G})$ $=\frac{1}{\alpha\beta\lfloor\frac{N}{n}\rfloor}\sum_{c=1}^{\alpha\lfloor\frac{N}{n}\rfloor}\sum_{i=1}^{\beta}\Big{[}\big{(}f^{\prime}(\boldsymbol{s}^{1,i}_{c},\mathcal{G}_{c}^{0})-\frac{1}{\beta}\sum_{j=1}^{\beta}f^{\prime}(\boldsymbol{s}^{1,j}_{c},\mathcal{G}_{c}^{0})\big{)}\nabla\ \log\pi_{c}(\boldsymbol{s}^{1,j}_{c}|\mathcal{G}_{c}^{0},\Omega_{c}^{0},\theta)\Big{]},$
$\nabla\mathcal{L}_{c2}(\mathcal{G})$ $=\frac{1}{\alpha\beta\lfloor\frac{N}{n}\rfloor}\sum_{c=1}^{\alpha\lfloor\frac{N}{n}\rfloor}\sum_{i=1}^{\beta}\Big{[}\big{(}f^{\prime}(\boldsymbol{s}^{2,i}_{c},\mathcal{G}_{c}^{1})-\frac{1}{\beta}\sum_{j=1}^{\beta}f^{\prime}(\boldsymbol{s}^{2,j}_{c},\mathcal{G}_{c}^{1})\big{)}\nabla\ \log\pi_{c}(\boldsymbol{s}^{2,j}_{c}|\mathcal{G}_{c}^{1},\Omega_{c}^{1},\theta)\Big{]},$
@@ 3.3 Application: Applicability in General CO Problems
This subsection discusses the applicability of UDC to various CO problems. Most existing neural divide-and-conquer methods rely on problem-specific designs, so they are only applicable to limited CO problems. UDC uses no heuristics algorithms or problem-specific designs in the whole solving process, thus can be applied to general CO problems that satisfy the following three conditions:
@@
For the second condition, the proposed UDC is disabled on complex CO problems whose solution cannot be guaranteed to be legal through an autoregressive solution process (e.g., TSPTW). For the third condition, on certain CO problems such as the MIS problem on dense graphs or job scheduling problems, the solution of sub-problems has already been uniquely determined, so the conquering stages become ineffective.
@@ 4 Experiment
To verify the applicability of UDC in the general CO problem, we evaluate the UDC on 10 combinatorial optimization problems that satisfy the conditions proposed in Section 3.3, including TSP, CVRP, KP, MIS, ATSP, Orienteering Problem (OP), PCTSP, Stochastic PCTSP (SPCTSP), Open VRP (OVRP), and min-max multi-agent TSP (min-max mTSP). Detailed formulations of these problems are provided in the Appendix B.
Implementation By employing lightweight networks for dividing policies, UDC can be directly trained on large-scale CO instances (e.g., TSP500 and TSP1,000). To learn more scale-independent features, we follow the varying-size training scheme [44, 45] in training and train only one UDC model for each CO problem. The Adam optimizer [46] is used for all involved CO problems and the detailed hyperparameters for each CO problem are provided in Appendix D.1. Most training tasks can be completed within 10 days on a single NVIDIA Tesla V100S GPU.
Baselines This paper compares the proposed UDC to advanced heuristic algorithms and existing neural solvers for large-scale CO problems, including classical solvers (LKH [47], EA4OP [48], OR-Tools [49]), single-stage SL-based sub-path solvers (BQ [14], LEHD [8]), single-stage RL-based constructive solvers (AM [50], POMO[5], ELG [6], ICAM [16], MDAM [51], et al.), heatmap-based solvers (DIMES [35], DIFUSCO [39], et al.) and neural divide-and-conquer methods (GLOP [13], TAM [12], H-TSP[20], et al.). The implementation details and settings of all the baseline methods are described in Appendix G.
Table 3: Experiment results on 500-node to 2,000-node OP and PCTSP and SPCTSP. The 500-scale, 100-scale, and 2,000-scale problems contain 128, 128, and 16 instances, respectively. “bs-$k$“ represents $k$-width beam search [50]. The overall best performance is in bold and the best learning-based method is marked by shade.
@@
Methods Obj.$\uparrow$ [^2] OP aims to maximize the objective function, which PCTSP and SPCTSP aim to minimize the objective function. Gap Time Obj.$\downarrow$ [^2] OP aims to maximize the objective function, which PCTSP and SPCTSP aim to minimize the objective function. Gap Time Obj.$\downarrow$ [^2] OP aims to maximize the objective function, which PCTSP and SPCTSP aim to minimize the objective function. Gap Time
@@
Results on 500-node to 2,000-node Instances. We first examine the performance of UDC models on large-scale CO problems ranging from 500-node to 2,000-node. Both single-stage SL-based constructive methods [16] and single-stage sub-path solvers [8] cannot be directly trained on this scale, so we report the results of models trained on the maximum available scale for these baselines. For UDC, we provide not only the greedy results (UDC-$\boldsymbol{x}_{2}$) but also the results after conducting more conquering stages (i.e., 50 and 250 stages, labeled as UDC-$\boldsymbol{x}_{50}$, UDC-$\boldsymbol{x}_{250}$). In the following conquering stages after a greedy solution, the starting point of sub-problem dividing is no longer fixed at $\frac{n}{2}$ but is sampled from Uniform$(0,n)$. For better performances, we sample $\alpha=50$ initial solutions $\boldsymbol{x}_{0}$ for TSP and CVRP. The comparison results of TSP and CVRP are shown in Table 2 and the results of OP, PCTSP, and SPCTSP are exhibited in Tabel 3, UDC demonstrates the general applicability, exhibiting consistent competitive performances in all the five involved CO problems. After multiple conquering stages, UDC can achieve significantly better results.
@@
UDC variants demonstrate the best results in PCTSP, SPCTSP, CVRP1,000, and CVRP2,000, together with the best learning-based method in most involved CO problems. The UDC-$\boldsymbol{x}_{250}$($\alpha$=1) showcases extensive efficiency in OP, PCTSP, and PCTSP. Compared to sub-path solvers BQ [14] and LEHD [8], UDC variants exhibit more competitive results on TSP1,000, TSP2,000, CVRP, and OP. RL-based constructive solver ICAM [16] with 8 augmentation [5] is outstanding performance on TSP500, TSP1,000, and CVRP500 (i.e., relatively smaller scale) but UDC variants demonstrates better effect and efficiency in other test sets. Moreover, compared to existing neural divide-and-conquer methods GLOP [13], H-TSP [20], and TAM(LKH3) [12], UDC achieves significant advantages in performance with no heuristic design, which preliminarily implies the significance of a unified training scheme for neural divide-and-conquer methods.
@@
Results on Benchmark Datasets. We also evaluate the proposed UDC on benchmark datasets TSPLib [52] and CVRPLib [53, 54]. We adopt the instances with 500-node to 5,000-node in TSPLib, CVRP Set-X [53], and very large-scale CVRP dataset Set-XXL as test sets. Instances in these datasets often with special distributions and can better represent real-world applications. As shown in Table 4, UDC variant UDC-$\boldsymbol{x}_{250}$ exhibits competitive benchmark results, demonstrating the best learning-based result on CVRPLib. The proposed UDC can also demonstrate superiority in the other 5 involved CO problems, very large-scale instances. Results of these experiments are shown in Appendix D.2 and Appendix D.5, respectively. And as to AppendixD.3 and AppendixD.4, UDC can also exhibit outstanding out-of-domain generalization ability on TSP, CVRP, and MIS instances with different distributions.
@@ 5 Discussion
The experimental results have preliminarily proven that the proposed UDC method achieves excellent results on a wide range of large-scale CO problems with outstanding efficiency. In this section, we conduct ablation experiments to verify the necessity of components in training UDC models. Appendix E includes the ablation experiments on the testing procedure of UDC, including the ablation on the number of conquering stages $r$ and the number of sampled solutions $\alpha$.
@@ 5.1 Unified Training Scheme versus Separate Training Scheme
For neural divide-and-conquer methods, the unified training process does not require special designs for pre-training conquering policies, thus being easier to implement compared to separate training processes. Moreover, unified training can also avoid the antagonism between the dividing and conquering policies. To validate this, we pre-train a conquering policy (i.e., ICAM) on uniform TSP100 data and then train two ablation variants of the original UDC (separate training scheme): one variant uses the pre-trained conquering policy in the subsequent joint training (i.e., Pre-train + Joint Training, similar to H-TSP [20]), and the other variant only train the dividing policy afterward (i.e., Pre-train + Train Dividing, similar to TAM [12] & GLOP [13]). According to the training curves of UDC and the two variants in Figure 3, the unified training scheme demonstrates superiority, while the Pre-train step leads the training into local optima, which significantly harms the convergence of UDC.
@@
Figure 3: Training curves of UDC variants with different training schemes.
Table 5: Results of UDC variants with various conquering policies.
@@
5.2 Ablation study: Conquering Policy
@@
In Appendix E.4, we further validate the conquering-policy-agnosticism of UDC on KP. We also conduct more other ablation studies to demonstrate the necessity of other components of the proposed UDC, including the Reunion step in the DCR training method and the coordinate transformation in the conquering stage.
@@ 6 Conclusion, Limitation, and Future Work
This paper develops a novel training method DCR for neural divide-and-conquer methods by alleviating the negative impact of sub-optimal dividing results. By using DCR in training, this paper enables a superior unified training scheme and proposes a novel unified neural divide-and-conquer framework UDC. UDC exhibits not only outstanding performances but also extensive applicability, achieving significant performance advantages in 10 representative large-scale CO problems.
Limitation and Future Work. Although UDC has achieved outstanding performance improvements, we believe that UDC can achieve better results through designing better loss functions. In the future, we will focus on a more suitable loss for the UDC framework to further improve training efficiency. Moreover, we will attempt to extend the applicability of UDC to CO problems mentioned in Section 3.3 that are not applicable in the current version.
@@
In the following Appendix, we will provide a literature review A, detailed definitions of CO problems & sub-CO problems B, details of the proposed UDC method C, additional experimental results D, additional ablation study E, and visualizations on a TSP1,000 instance, a CVRP1,000 instance (i.e., X-n1001-k43 in CVRPLib), and a OP1,000 instance F.